OpenStudy (anonymous):

A 20 kg block of ice slides without friction down a long hill. The start of the hill is 7.5m higher than the bottom of the hill, and the path traveled is 18.0m. Find the speed of the block when it reaches the bottom of the hill.

OpenStudy (anonymous):

\[v ^{2}=u ^{^{2}}+2as\] a=mgsin(theta) \[\sin \theta =\frac{ 7.5 }{ 18 }\]

OpenStudy (anonymous):

u=0 54.22